Happy Easter All, I am having a problem writing an equation containing a given point and parallel. (4,9) x+4y=5, do you think you could help me?
Answer by Alan3354(69443)

given point and parallel. (4,9) x+4y=5
-----------------
Find the slope of the line. To find the slope, m, put it into slope-intercept form, y = mx + b. That means solve for y.
x+4y=5
4y = -x + 5
y = (-1/4)x + 5/4
m = -1/4
All lines parallel to this one will have the same slope.
---------------
then use
y-y1 = m*(x-x1) where (x1,y1) is the point.
y-9 = (-1/4)*(x-4)
y-9 = (-1/4)x + 1
y = (-1/4)x + 10 slope-intercept form
x + 4y = 40 standard form